# Sampling-based Distributed Training with Message Passing Neural Network

### v2

@@ 5 Ablation Studies on Scalability and Communication Overhead
We also increase the sampled node count $s$ to ascertain the impact of scalability and communication overhead. This increase in node samples becomes imperative for larger numerical grids. Figure 7 shows the increase in the training speed of DS-MPNN4 compared to S-MPNN. The increase in speed is attributed to the reduction in the time required for graph kernel $\mathcal{G}$ generation within DS-MPNN4, as a consequence of decreased number of edge formations required across each GPU domain. Correspondingly, the percentage of training time spent on communication decreases as the node count increases, because, while the aggregate data for communication grows with the node count, the substantial bandwidth of GPUs maintains a constant communication time.
@@
Figure 8: Training and inference time as a function of number of distributed domains.
@@
6 Conclusions
We address the issue of scaling edge-based graph methods to larger numbers of nodes by introducing distributed training for message-passing neural networks (MPNNs). When combined with node sampling techniques, this distributed approach allows us to scale a larger number of nodes with losing no or minimal loss in accuracy as compared to the single-GPU implementation while achieving a considerable decrease in training and inference times. We also show comparisons to the graph convolution networks (GCNs) and establish that edge-based methods like MPNNs outperform node-based methods like GCNs - This work opens up new avenues for the use of edge-based graph neural networks in problems of practical interest where the number of nodes can be impractically large, $\sim O(10^{5}-10^{6})$, for single-GPU memory limits.  The deployment of DS-MPNN is complex, notably in managing unstructured grids and their adjacent elements. However, the domain of distributed graph training is advancing swiftly. Notably, libraries tailored for these challenges have been introduced into the latest iteration of PyTorch Geometric (Fey & Lenssen, 2019), version 2.5.0.
7 Impact statements
This paper presents work whose goal is to advance the field of graph neural networks applied to physical systems. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Training Algorithm
Algorithm 1 Training of DS-MPNN.

### v3

@@ Appendix A Ablation Studies on Scalability and Communication Overhead
Figure 8: Training and inference time as a function of the number of distributed domains.
@@
Appendix B Darcy Flow Equations
2-D Darcy flow equation on the unit box is a second-order elliptic PDE given as:
$-\nabla(a(x)\nabla u(x))=f(x)\quad\forall x\in(0,1)^{2}$ (4)
$u(x)=0\quad\forall x\in\partial(0,1)^{2}.$ (5)
Here, $a$ and $f$ are the spatially varying diffusion coefficients and the forcing field, respectively. $u$ is the solution field on the 2-D domain. This example parallels the approach used in the Graph Neural Operator (GNO) as discussed by  Li et al. (2020a), adapted for a single GPU setup with a notable distinction: we update the edge attributes following each radius hop or message-passing step, and both edges and nodes are subjected to sampling. The experiment investigates the mapping $a\rightarrow u$.
Appendix C Darcy Flow Visualizations
Figure 9 illustrates that both DS-MPNN and S-MPNN configurations provide similar predictions for the Darcy flow dataset. In contrast, the GCN model significantly under-performs, as it excessively smooths the solution and fails to capture the roughness at the boundaries accurately.
Appendix D AirfRANS equations
.$\nabla.\mathbf{U}=0.$ (7)
Appendix E Three-dimensional step flow visualizations
The predictions for S-MPNN and DS-MPNN4 are further illustrated in Figure 10, where two slices are presented: one at $z=0.5$ in the xy-plane, and the other at $x=3.5$, which lies within the recirculation region. The results demonstrate the accuracy and consistency of both models. The scatter plot in Figure 11 further corroborates this by comparing the predictions at $z=0.5$ and $x=3.5$, indicating a high degree of agreement in the predictions of two models.
Figure 10: Streamwise $u$-velocity prediction (left) and wall-normal $v$-velocity (right). The xy-slice is at $z=0.5$ and the yz-slice is at $x=3.5$. The top row depicts S-MPNN predictions, while the bottom row shows DS-MPNN4 predictions.
Figure 11: Comparison of S-MPNN and DS-MPNN4 at $x=3.5$ and $z=0.5$.
Appendix F Training Algorithm